OpenStudy (anonymous):

The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1150. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many one-step equations auditioned today? 430 720 1150 1300

hero (hero):

lol, thanks

hero (hero):

x = one step y = two steps

hero (hero):

x + y = 1150 3x - 2y = 1300

hero (hero):

y = 1150 - x y = (3x - 1300)/2 y = y 2(1150 - x) = 3x - 1300 2300 - 2x = 3x - 1300 2300 + 1300 = 5x 3600 = 5x x = 3600/5 x = 720 y = 430

hero (hero):

720 one-step equations auditioned today